# Identifying and Addressing Delusions for Target-Directed Decision Making

### v2

@@ 1 Introduction
In psychiatry, delusions depict the inability to reject false beliefs (Kiran & Chaudhury, 2009), which in the RL context translates to false beliefs that are natural results of an agent’s learning. Particularly problematic for target-directed agents, when false beliefs cannot be resolved through training, delusional behaviors may appear in the form of chasing nonexistent or unreachable targets, or those beyond safety constraints (Bengio et al., 2024). Delusions limit the agents’ situational understanding, and lead to even worse generalization in the target tasks. Delusions stem from the incoordination of the generator and the estimator, respectively: problematic targets should not be generated, and even if they were, they should not be favored(Corlett, 2019). While delusions have various sources, this paper focuses on the misevaluation by the estimator caused by improper training processes, a combination of inadequate update rules and inappropriate training data distributions.
@@ 3.1 Generator Delusions
The second type includes those that correspond to states in the task MDP, but (some) cannot be reached from the current state. Unlike G.3.1, these targets are only temporarily inappropriate. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$. Naïvely applying relabeling strategies, e.g.episode”, to generator training, can be a major cause of G.3.1.1 hallucinations. Similarly, G.3.1.1 may not result in delusional behaviors if the matching Type E.3.2.2 delusions are not produced by the estimator.
@@ 3.2 Estimator Delusions
An estimator evaluates the favorability of proposed targets, based on some relationship between the current state and the candidates, which often involve assessing the implementability of targets, represented as, e.g., distances, etc. Acting as a firewall, the estimator of a target-directed agent can assist the selection among the proposed targets, bearing the potential to filter out delusional behaviors. Agents without estimators, e.g. Hafner et al. (2022), must accept proposed targets unconditionally, and thus are at significant risk of delusions (chasing G.3.1 and G.3.1.1). Importantly, the incorporation of estimators merely delegates the responsibility of addressing delusional behaviors to the estimator.
Estimators’ delusional beliefs about targets can directly cause delusional behaviors. For an estimator to learn to reject delusions, we need 1) effective learning rules that can punish the favorability of problematic targets, and 2) training data that can counteract delusions through the learning rules. A lack of either can result in blind spots of knowledge about targets, delusional or non-delusional.
@@
Training Data Distribution Even with capable update rules, delusions can arise from training data distributions that lack the exposure to the data counteracting false beliefs in the estimator. For example, some training data distributions suffer from low-diversity, e.g., training target-directed agents relabeling mechanisms (Dai et al., 2021). But relabeling isn’t necessarily the solution: 1) Certain relabeling strategies intrinsically cause delusions. For instance, “future” only relabels future observations, and thus only exposes a learner to reachable targets, confusing the estimator when a “past” target is generated at decision time; 2) Trajectory-level relabeling is also problematic. Short trajectories, common in many training procedures, prevent estimators from learning to evaluate distant targets. Procedures like “episode” limit exposure, risking delusions. Short trajectories can be a product of experimental design (initial state distributions, episode lengths) or innate environmental characteristics (density of terminal states). For example, to train useful skills in a hard-exploration problem, a popular technique is to relax the initial state distribution of the training tasks, thus producing shorter but more efficient trajectories (Erraqabi et al., 2021); In some other cases, with dense terminal states or short Maximum Episode Lengths (MELs), each experienced trajectory generally only covers a limited portion of the state space.
@@ 4.1 Assistive Strategies for Addressing Delusions
The following strategies seek to expand the support of the training data distribution, to include those source-target combinations involving G.3.1 and G.3.1.1 targets, respectively. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
@@ 6.3 Methods
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence, among all, is chosen to be the target and used to condition a lower-level policy. LEAP is prone to delusions because its planned sequences require all subgoals to be non-delusional (Nasiriany et al., 2019).
@@ 6.6 Summary of Experiments & Empirical Guidelines
6.7 Empirical Guidelines
Training generators with problematic targets could cause hallucinations, which increases risks for E.3.2.1 and E.3.2.2 induced delusions; Incorporate an estimator to enable rejection of unwanted targets; Use update rules with appropriate properties and try to maximize the diversity of the training data distribution (against E.3.2); Inspect the candidate targets if possible, since their low quality may indicate higher E.3.2.1 risks, which we can mitigate with the idea behind “generate”; Analyze the state structure of the target tasks and identify temporary unreachabilities. These E.3.2.2 risks may indicate the effectiveness of the idea behind “pertask”.
@@ 7 Conclusion & Future Work
We discussed the causes and mitigation of delusions, a shared failure mode of target-directed agents. When applying our strategies on HER-trained agents, the resulting hybrid hindsight relabeling approaches flexibly satisfy the needs of the generator and the estimator, while avoiding the naïveté in training processes that could lead to more delusional behaviors.
@@ 8 Reproducibility Statement
The results presented in the experiments are fully-reproducible with the source code published at https://github.com/pwnerharry/delusions.
9 Acknowledgements
Mingde is grateful for the financial support of the FRQ (Fonds de recherche du Québec) and the collaborative efforts during his internship at Borealis AI, Montreal. We appreciate the computational resources allocated to us from Mila and McGill university.

### v3

@@ 1 Introduction
Delusions are obviously wrong beliefs that some agents hold despite contradictory evidence. Formally, they reflect the inability to reject false beliefs (Kiran & Chaudhury, 2009), which in the RL context translates to false beliefs that are natural results of an agent’s improper learning process. Particularly problematic for target-directed agents, delusional behaviors appear in the form of chasing nonexistent or unreachable targets, or those beyond safety constraints (Bengio et al., 2024), which in-turn limit the agents’ situational understanding, and lead to worse generalization in the target tasks or even safety catastrophes. Delusions stem from the incoordination of the generator and the estimator, respectively: problematic targets should not be generated, and even if they were, they should not be favored (Corlett, 2019). While delusions can have various sources, this paper focuses on the misevaluation by the estimator caused by improper training processes, a combination of inadequate update rules and inappropriate training data distributions.
@@ 3.1 Generator Delusions
The second type includes those that correspond to states in the task MDP, but cannot be fulfilled from the current state. Unlike G.3.1, these targets are only temporarily inappropriate. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$. Naïvely applying relabeling strategies, e.g.episode”, to generator training, can be a major cause of G.3.1.1 hallucinations. Similarly, G.3.1.1 may not result in delusional behaviors if the matching Type E.3.2.2 delusions are not produced by the estimator.
@@ 3.2 Estimator Delusions
An estimator evaluates the favorability of proposed targets, based on some relationship between the current state and the targets, which often involve assessing the implementability of targets, represented as, e.g., distances, etc. Acting as a firewall, the estimator of a target-directed agent can enable the selection among the proposed targets, bearing the potential to filter out hallucinated targets. Agents without estimators, e.g. Hafner et al. (2022), must accept proposed targets unconditionally, and thus are at significant risk of chasing problematic targets.
Incorporating an estimator does not mean delusions addressed, as its delusional beliefs about targets can directly cause delusional behaviors. For an estimator to learn to reject delusions, we need 1) effective learning rules that can punish the favorability of problematic targets, and 2) training data that can counteract delusions through the learning rules. A lack of either can result in blind spots of knowledge about targets, delusional or non-delusional.
@@
Training Data Distribution Even with capable update rules, delusions can still arise from the lack of exposure to data counteracting the estimators’ false beliefs. As shown in Fig. 1, training data are created by organizing mechanisms based on experienced trajectories, e.g. hindsight relabeling for HER-trained agents. Despite these mechanisms can lead to more diversity in training data, they are often not properly set to counter delusions. Take hindsight relabeling as an example: 1) Certain relabeling strategies intrinsically cause delusions. For instance, “future” only relabels future observations, thus only exposes a learner to reachable targets, confusing the estimator when a “past” target is evaluated at decision-time; 2) Trajectory-level relabeling can also be problematic. Short trajectories, common in many training procedures, only cover limited portions of the state space and prevent estimators from learning to evaluate distant targets. Short trajectories can be a product of experimental design (initial state distributions, Maximum Episode Lengths, MELs (Erraqabi et al., 2021)) or innate environmental characteristics (density of terminal states). It is worth noting that delusions caused by improper training data (organizing mechanisms) also extend to cases like few-shot generalization.
@@ 4.1 Assistive Strategies for Addressing Delusions
The following strategies, which creates the training data from experienced episodes, seek to expand the support of the training data distribution, to include those source-target combinations involving G.3.1 and G.3.1.1 targets, respectively. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
@@ 6.3 Methods
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence is chosen to be the target, and used to condition a lower-level policy. LEAP is prone to delusions, since one problematic subgoal and destroy a whole sequence (Nasiriany et al., 2019).
@@ 6.6 Summary of Experiments & Empirical Guidelines
7 Empirical Guidelines
The following steps can be applied to target-directed methods beyond HER:
  1. 1.
    Incorporate an estimator to enable rejection of unwanted targets;
  2. 2.
    Use update rules with appropriate properties and try to maximize the diversity of the training data distribution (against E.3.2);
  3. 3.
    Inspect the candidate targets if possible, since their low quality may indicate higher E.3.2.1 risks, which we can mitigate with the idea behind “generate”;
  4. 4.
    Analyze the state structure of the target tasks and identify temporary unreachabilities. These E.3.2.2 risks may indicate the effectiveness of the idea behind “pertask”.
@@ 8 Conclusion & Future Work
We identified the types and the causes of delusional behaviors, a shared failure mode of target-directed agents. When applying our strategies on HER, the resulting hybrid hindsight relabeling strategies flexibly satisfy the needs of the generator and the estimator, while avoiding the naïveté in training processes that could lead to more delusional behaviors and catastrophic safety breaches.
@@ 9 Reproducibility Statement
The results presented in the experiments are fully-reproducible with the submitted source code.